OpenStudy (anonymous):

If the Federal Reserve decreases the reserve rate from 5% to 2%, how does this affect the amount of money that would result because of fractional reserve banking from an initial deposit into a bank of $25,000? A. It increases the amount by $500,000. B. It decreases the amount by $750,000. C. It increases the amount by $750,000. D. It decreases the amount by $500,000.

OpenStudy (anonymous):

With a 5% reserve ratio, the Fed can create up to \(\dfrac{1}{.05}=20\) times the initial deposit. With a 2% reserve ratio, it can create up to \(\dfrac{1}{.02}=50\) times the initial deposit. So \(lowering\) the reserve ratio would \(increase\) the amount of money made. \(50x-20x=30x\), where \(x=$25,000\).